1. What is the minimum value of the function y = 3x2 + 2x – 8?

2. The equation h = 40t – 16t2 describes the height h, in feet, of a ball that is thrown straight up as a function of the time t, in seconds, that the ball has been in the air. At what time does the ball reach its maximum height? What is the maximum height?

Answer by oscargut(2103)

Vertex is V=(-b/2a,f(-b/2a))
1)f(x)=3x^2+2x-8 minimum is at x=-2/6=-1/3 f(-1/3)=-25/3 (minimum value)
2)f(t)= – 16t^2+40t maximum is at -40/-32=5/4, f(5/4)=25
t= 5/4 sec height is 25 feet